# Multimodal machine learning with large language embedding model for polymer property prediction

### v1

@@ 3 Discussions
Table 2: Comparison of predictive performance (mean $R^{2}$ scores ± standard deviation) across various polymer properties for different models. Results were obtained using five-fold cross-validation on test datasets. PolyLLMem refers to the multimodal model integrating LLM-generated text embeddings and Uni-Mol structural embeddings. LLM+XX and Uni-Mol+XX denote models utilizing embeddings from Llama 3 or Uni-Mol, respectively, as input features for the indicated methods (MLP, XGB). MF+XGB and Descriptors+XGB denote models using RDKit molecular fingerprints (MF) or RDKit molecular descriptors as input features for XGB. Best-performing results per property are highlighted in bold with arrows.
@@
After confirming that the LLM embeddings using Llama 3 retained essential chemical information, we evaluated their predictive performance on over 22 distinct polymer properties. Initially, the generated 4096-dimensional embeddings for each polymer were used as input to a XGBoost model (LLM+XGB)52. The performance results for various polymer properties were summarized in Table 1, which reports the average $R^{2}$ values obtained from five-fold cross-validation on the test set. LLM+XGB achieved good performance on several polymer properties, with $T_{g}$, $E_{gc}$, $E_{gb}$, $\mu_{H_{2}}$, each exhibiting $R^{2}$ value above 0.8. These results support our UMAP analysis, indicating that the LLM embeddings capture meaningful chemical information and perform well on certain properties prediction tasks. However, when compared with benchmark models trained on MF features (MF+XGB) or molecular descriptors (descriptors+XGB), LLM+XGB consistently underperformed across most property predictions, particularly when contrasted with MF+XGB. In addition, we evaluated other ML models such as RF, Ridge Regression, AdaBoost, MLP, etc. using the LLM-generated embeddings. Among these, the model combining LLM embeddings with an MLP (LLM+MLP) demonstrated exceptional performance, yielding results that were comparable to those of the benchmark models on the majority of the property prediction tasks, as shown in Table 1. Additional results for other models were provided in the SI, as their performances were inferior to those of the benchmark models. We also trained MLP models using MF features and molecular descriptors respectively, but these results were inferior to those obtained with MF+XGB or descriptors+XGB (details were provided in the SI). Moreover, Uni-Mol embeddings alone as input features for ML models were shown in Table 1 as comparison.
@@
Although LLM+MLP demonstrated exceptional performance overall, there remains room for improvement since the gains in prediction accuracy were modest and some property prediction tasks still underperformed compared to the baseline models. LLM based embeddings primarily capture textual information from PSMILES strings, leaving out crucial aspects of a molecule’s structure. To address this gap, we introduced Uni-Mol, a deep learning architecture that encodes the 3D geometry and conformational characteristics of small molecules, to generate embeddings that capture structural information, and integrated them with LLM-based embeddings to form the multimodal model PolyLLMem. As shown in Table 1, PolyLLMem demonstrated superior performance on most property prediction tasks compared to the baseline models, except for certain mechanical properties( $\sigma_{y}$, $\epsilon_{b}$), crystallization tendency $X_{c}$ and conductivity $\sigma$. Additionally, we compared the prediction results of PolyLLMem with those of other graph-based and transformer-based models, such as PolymerBERT, TransPolymer, and PolyGNN, as detailed in the SI. Given the limited training data used in our study, PolyLLMem exhibits performance that is comparable to, and in some cases exceeds, that of these benchmark models on the majority of polymer properties. However, PolyLLMem’s performance in predicting gas permeability was not as competitive as that of the graph-based and transformer-based models. Several factors may contribute to this discrepancy. One potential factor is that the training dataset may have introduced differences. We used the latest experimental gas permeability data from Phan et al. 47, which is more diverse than the data used to train PolymerBERT and PolyGNN in their original studies. In fact, even for a single-task prediction, PolyGNN’s performance with this new dataset in their latest work 47 was lower than reported in its original paper 28. Moreover, PolymerBERT employed multitask training, which may enable it to capture a broader range of information compared to our single-task approach, resulting in improved performance for certain properties. Additionally, our findings indicate that PolyLLMem struggled to accurately predict mechanical properties, suggesting that further model refinement or additional data may be necessary to enhance its performance in this area.
@@
Figure 4: Token-level attribution analysis for $T_{g}$ prediction using PolyLLMem of selected polymers: (a) [*]CC([*])C, (b) [*]CC([*])c1ccncc1, (c) [*]CC([*])(F)C(=O)OCC(F)(F)C(F)(F)F. Token level attributions were computed using Integrated Gradients, highlighting the contribution of individual chemical tokens to model predictions. High attribution values indicate greater significance in determining the predicted property. Red circles are used to label the atoms that have the highest attributions
@@
Lastly, we evaluated the interpretability of the model. In our model, we extracted the token-level embeddings and applied a mean pooling layer to obtain the aggregated representation used by PolyLLMem. Using this setup, we computed Integrated Gradients 53 along the path from a zero (baseline) input to the actual token-level embeddings by passing them through a wrapper function that performs the mean pooling, thereby generating token-level attributions. By analyzing the aggregated attribution values for each token, we identified those that contributed most significantly to the predictions. Figure 4 shows the token attributions for $T_{g}$ prediction from the selected polymers. For polymer [*]CC([*])C in Figure 4a, the tokens representing the carbon backbone (CC) and the polymerization point tokens ([*]) show relatively high aggregated attribution values, indicating that the model emphasizes the alkyl backbone and its connection points. In Figure 4b, for polymer [*]CC([*])c1ccncc1, the tokens corresponding to the nitrogen containing ring (n) as highlighted by the red marker, and the ring closure marker (1) exhibit higher aggregated attribution values, suggesting that the model focuses on these structural features. In polymer chemistry, heteroaromatic rings, particularly those containing nitrogen, can significantly affect chain packing and polarity, which in turn affects $T_{g}$. Lastly, in polymer [*]CC([*])(F)C(=O)OCC(F)(F)C(F)(F)F, which exhibits one of the highest $T_{g}$ values in our dataset, the tokens corresponding to the trifluoromethyl (–CF3) groups have the highest aggregated attribution values, suggesting the model strongly focuses on these fluorinated groups. This also aligns with real world polymer chemistry, where fluorinated groups can significantly affect chain rigidity and consequently increase $T_{g}$. Collectively, the model’s attention to these functional groups and structural moieties demonstrates its ability to capture the key chemical features that determine $T_{g}$ in practice.
@@ 4 Conclusion
In this study, we introduced a simple yet effective multimodal framework, PolyLLMem, which synergistically integrates LLM-based text embeddings with 3D molecular structure embeddings from Uni-Mol to predict a wide range of polymer properties. By leveraging PSMILES strings as a source of rich textual information and integrating them with geometrical descriptors, our approach successfully encapsulates both the chemical context inherent in polymer notations and the critical conformational features of polymer molecules. Our extensive evaluation across 22 distinct property prediction tasks demonstrated that PolyLLMem achieves competitive performance compared to established transformer and graph-based models that are specifically designed for the polymer domain. Remarkably, while those models, especially transformer-based models, typically require training on millions of data points, our model was developed using a relatively small dataset, underscoring its simplicity and effectiveness. Notably, the LLM embeddings model alone (LLM+XGB) showed promise for properties such as $T_{g}$, $E_{gc}$, and $E_{gb}$, indicating some of the chemical information within polymer domains already embedded in the LLM. To integrate the LLM embeddings model with 3D structural information further enhanced predictive accuracy across most tasks. Additionally, using Integrated Gradients also revealed that the model effectively identifies key chemical motifs and structural features in line with established chemical intuition.
Despite these promising results, challenges remain, particularly in predicting certain mechanical properties. These discrepancies highlight the inherent complexities of polymer data and indicate the need for further refinement of data augmentation techniques and model architectures. Future work could enhance the embedding model’s complexity by leveraging token-level embeddings in conjunction with multi-head attention mechanisms, which may further improve the accuracy of polymer property predictions. Overall, our work underscores the potential of combining LLM embeddings with molecular structure information to accelerate the discovery and optimization of polymer materials.
@@ 4.1 Code availability
The PolyLLMem code is available for academic use at https://github.com/zhangtr10/PolyLLMem.

### v2

@@ 3.0.2 Performance of PolyLLMem
Table 2: Comparison of predictive performance (mean $R^{2}$ scores ± standard deviation) across various polymer properties for different models. Results were obtained using five-fold cross-validation on test datasets. PolyLLMem refers to the multimodal model integrating LLM-generated text embeddings and Uni-Mol structural embeddings. LLM+XX and Uni-Mol+XX denote models utilizing embeddings from Llama3 or Uni-Mol, respectively, as input features for the indicated methods (MLP, XGB). MF+XGB and Descriptors+XGB denote models using RDKit molecular fingerprints (MF) or RDKit molecular descriptors as input features for XGB. Best-performing results per property are highlighted in bold with arrows. The properties of gas permeabilities ($\mu_{x}$), tensile strength at break ($\sigma_{b}$), tensile strength at yield ($\sigma_{y}$), tensile strength at yield ($\sigma_{y}$), elongation at break ($\epsilon_{b}$), Young’s modulus($E$) and conductivity ($\sigma$) were trained on log scale and the $R^{2}$ value for those properties were reported on this scale.
Figure 4: Scatter plots of ground truth vs. predicted values for the selected properties: (a) $T_{g}$, (b) $\rho$, (c) $T_{d}$, (d) $T_{m}$, (e) $E_{gc}$, (f) $E_{at}$, (g) $E$, (h) $\mu_{CH_{4}}$, (i) $\mu_{O_{2}}$. The $R^{2}$ value for properties of $E$, $\mu_{CH_{4}}$, $\mu_{O_{2}}$ were calculated based on the training in the log scale, whereas the MAE were reported on the original value.
@@
After confirming that the LLM embeddings using Llama3 retained essential chemical information, we evaluated their predictive performance on over 22 distinct polymer properties. Initially, the generated 4096-dimensional embeddings for each polymer were used as input to a simple XGBoost model (LLM+XGB)52. The performance results for various polymer properties were summarized in Table 1, which reports the average $R^{2}$ values obtained from five-fold cross-validation on the test set. Despite straightforward, LLM+XGB already achieved satisfying performance on several polymer properties, including $T_{g}$, $E_{gc}$, $E_{gb}$, $\mu_{H_{2}}$, each exhibiting $R^{2}$ value above 0.8. These results support our UMAP analysis, indicating that the LLM embeddings capture meaningful chemical information and perform well on certain property prediction tasks. However, when compared with benchmark models trained on MF features (MF+XGB) or molecular descriptors (descriptors+XGB), LLM+XGB consistently underperformed across most property predictions, particularly when contrasted with MF+XGB. In addition, we evaluated other ML models such as RF, RR, AdaBoost, MLP, etc. using the LLM-generated embeddings. Among these, the model combining LLM embeddings with an MLP (LLM+MLP) demonstrated exceptional performance, yielding results that were comparable to those of the benchmark models on the majority of the property prediction tasks, as shown in Table 1. Additional results for other models were provided in the SI, as their performances were inferior to those of the benchmark models. We also trained MLP models using MF features and molecular descriptors respectively, but these results were inferior to those obtained with MF+XGB or descriptors+XGB (details were provided in the SI).
@@
Although the combination of LLM+MLP alone demonstrated exceptional performance overall, there remains room for improvement since the gains in prediction accuracy were modest and some property prediction tasks still underperformed compared to the baseline models. As LLM-based embeddings primarily capture textual information from PSMILES strings, crucial aspects of a molecule’s structure might be left out . To address this gap, we introduced Uni-Mol, a deep learning architecture that encodes the 3D geometry and conformational characteristics of small molecules, to generate embeddings that capture structural information. Integrating Uni-Mol-based embeddings with LLM-based ones, our multimodal model PolyLLMem was formed. As shown in Table 1 and Figure 4 , PolyLLMem demonstrated superior performance on most property prediction tasks compared to the baseline models, except for certain mechanical properties ($\sigma_{y}$, $\epsilon_{b}$), crystallization tendency $X_{c}$ and conductivity $\sigma$. (Note that Uni-Mol embeddings alone as input features for ML models were also shown in Table 1 as a comparison.) Additional comparisons of prediction results of PolyLLMem with those of other state-of-the-art graph-based and transformer-based models, such as PolymerBERT, TransPolymer, and PolyGNN, as detailed in the SI. PolyLLMem exhibits performance that is comparable to, and in some cases exceeds, that of these benchmark models on the majority of polymer properties. Given the limited training data used in our study, this highlights the model’s strong data efficiency and generalization capability.
3.0.3 Token-Level Interpretability
@@
Figure 5: Token-level attribution analysis for $T_{g}$ prediction using PolyLLMem of selected polymers: (a) [*]CC([*])C, (b) [*]CC([*])c1ccncc1, (c) [*]CC([*])(F)C(=O)OCC(F)(F)C(F)(F)F. Token-level attributions were computed using Integrated Gradients, highlighting the contribution of individual chemical tokens to model predictions. All values are normalized by the attribution value of token [*] for clarity. Higher attribution values indicate greater significance in determining the predicted property. Light purple and light blue are used to tightly the top 2 influential tokens, with arrows of the same color notes for the corresponding substructure. The backbone of polymers is shaded in lightgrey.
@@
Lastly, we commented on the interpretability of the model. The token-level embeddings were extracted using Llama3 tokenizer, following by appling a mean pooling layer to obtain the aggregated representation used by PolyLLMem. Using this setup, the Integrated Gradients was computed 53 along the path from a zero (baseline) input to the actual token-level embeddings. A wrapper function applied mean pooling over the embeddings, enabling the attribution scores to be assigned at the token level. Such value can be correlated to the token’s contribution to property prediction. As noted in the previous section regarding challenges with accurately splitting chemical notations, once we obtained the attribution values for each token, we applied the same merging strategy described earlier to consolidate and refine these values according to the refined token representations. Figure 5 shows the refined token attributions for $T_{g}$ prediction from the selected polymers. For polymer [*]CC([*])C in Figure 5a, the tokens representing the carbon backbone (CC) and the sidechain carbon (C) show relatively high aggregated attribution values, indicating that the model emphasizes the alkyl nature. In Figure 5b, for polymer [*]CC([*])c1ccncc1, the tokens corresponding to the nitrogen-containing ring (n), and the ring closure marker (1) exhibit higher aggregated attribution values, showing the importance of the substituent effect in the benzene. This is as expected, as in polymer chemistry, heteroaromatic rings, particularly those containing nitrogen, can significantly affect chain packing and polarity, which in turn affects $T_{g}$. Lastly, in polymer [*]CC([*])(F)C(=O)OCC(F)(F)C(F)(F)F, which exhibits one of the highest $T_{g}$ values in our dataset, the tokens corresponding to the fluorine in the fluoromethyl (including both –CF3 and -CF2-) groups have the highest aggregated attribution values. This also aligns with real-world polymer chemistry, where fluorinated groups can significantly affect chain rigidity and consequently increase $T_{g}$. Collectively, the model’s attention to these functional groups and structural moieties demonstrates its ability to capture the key chemical features that determine $T_{g}$ in practice.
@@ 4 Conclusion
In this study, we introduced a simple, lightweight, yet effective multimodal framework, PolyLLMem, which synergistically integrates LLM-based text embeddings with 3D molecular structure embeddings from Uni-Mol to predict a wide range of polymer properties. By leveraging PSMILES strings as a source of rich textual information and integrating them with geometrical descriptors, our approach successfully encapsulates both the chemical context inherent in polymer notations and the critical conformational features of polymer molecules. Our extensive evaluation across 22 distinct property prediction tasks demonstrated that PolyLLMem achieves competitive performance compared to established transformer and graph-based models that are specifically designed for the polymer domain without the requirements of millions of data points or data augmentation. Notably, the LLM (Llama3) embeddings model alone (LLM+XGB) showed promise for properties such as $T_{g}$, $E_{gc}$, and $E_{gb}$, indicating some of the chemical information within polymer domains was already embedded in the Llama3. Further integration of the LLM embeddings model with 3D structural information enhanced predictive accuracy across most tasks. Additionally, using Integrated Gradients also revealed that the model effectively identifies key chemical motifs and structural features in line with established chemical intuition.
Despite these promising results, challenges remain, particularly in predicting certain mechanical properties. These discrepancies highlight the inherent complexities of polymer data and indicate the need for further refinement of data augmentation techniques and model architectures. Future work shall focus on enhancing the embedding model’s complexity by leveraging token-level embeddings in conjunction with multi-head attention mechanisms, which may further improve the accuracy of polymer property predictions. Nonetheless, our work here already underscores the potential of combining LLM embeddings with molecular structure information to accelerate the discovery and optimization of polymer materials.
@@ 4.1 Code availability
Source code is available for academic use at https://github.com/zhangtr10/PolyLLMem.